OpenStudy (rsadhvika):

Show that days with the identical calendar date in the years 1999 and 1915 fell on the same day of the week.

OpenStudy (rsadhvika):

http://gyazo.com/7b69acf92895f6f87e047c6f7955c037

OpenStudy (ikram002p):

whats so special about 1999 :O or 1915

OpenStudy (ikram002p):

http://prntscr.com/58jsot

OpenStudy (ikram002p):

so basically we need to avoid one years each 4 years and count more 7 to get a year with same identical calendar date ... that would be each two 11 years and one 6 years

OpenStudy (ikram002p):

so, 1999-1915=0 mod 7 xD

OpenStudy (ikram002p):

bhahaha xD

OpenStudy (ikram002p):

this is nice to show what day next year would be

OpenStudy (rsadhvika):

I see... if you have time could you help me with solution

OpenStudy (ikram002p):

solution for what :O

OpenStudy (rsadhvika):

i mean proof

OpenStudy (ikram002p):

like proper proof ? ok , main i dea is to show that 1999-1915 =0 mod 7

OpenStudy (rsadhvika):

\[Y_1 \equiv Y2 \pmod {7}\] doesnt necessarily mean they will have exact same "day name" on each day of the entire year rigt ?

OpenStudy (rsadhvika):

If thats the case we will be having that pattern every 7 years i guess

OpenStudy (ikram002p):

nope , count 7+4=11 when we have leap in the start and 7-1 when we dont have

OpenStudy (rsadhvika):

thats for advancing a year ?

OpenStudy (ikram002p):

?